OpenStudy (anonymous):

Convert the Cartesian equation x^2 + y^2 - 6x = 0 to a polar equation.

OpenStudy (anonymous):

so, what are the polar forms of "x" and "y"?

OpenStudy (anonymous):

\[x=r\cos\theta\\y=r\sin\theta\]

OpenStudy (anonymous):

now, plug these for "x" and "y" in the given equation.

OpenStudy (anonymous):

Okay, hold on. & Thank you!

OpenStudy (anonymous):

Well obviously you get (rcostheta)^2 + (rsintheta)^2 - 6(rcostheta)=0 but I don't really know where to go from there. But I have to get r alone, right?

OpenStudy (anonymous):

correct. so you get \[r^2\cos^2\theta+r^2\sin^2\theta-6r\cos\theta=0\] factor out \(r^2\) from the first two terms!

OpenStudy (anonymous):

To factor it out it would give me r^2(cos^2theta + sin^2theta) - 6rcostheta = 0 Right?

OpenStudy (anonymous):

keep going.. see any Identities you could use there?

OpenStudy (anonymous):

\[\sin^2\theta+\cos^2\theta=?\]

OpenStudy (anonymous):

Doesn't that equal 1?

OpenStudy (anonymous):

exacly.

OpenStudy (anonymous):

use the equation button below to enter the next step that you have

OpenStudy (anonymous):

\[r^2 - 6xcos^2\theta = 0\] & I didn't even know that button existed... Sorry about that

OpenStudy (anonymous):

then it would be \[r^2=6rcos^2\theta\]

OpenStudy (anonymous):

Then you would take the square roots, right?

OpenStudy (anonymous):

well, we try not to do that second step.

OpenStudy (anonymous):

Oh so would the answer be \[r^2= 6\cos \theta\]

OpenStudy (anonymous):

\[r^2-6r\cos\theta=0\\ r(r-6\cos\theta)=0\\ r=0\qquad{\rm OR}\qquad r-6\cos\theta=0\implies\Large\boxed{r=6\cos\theta}\]

OpenStudy (anonymous):

the first one is rejecected since it is the "origin" and not a curve.

OpenStudy (anonymous):

Ohhh. I see. Thank you for your help :)